OpenStudy (anonymous):

write the equation of a line given the following information: passes through the points (-4,-5) and (6,0)

OpenStudy (anonymous):

To write the equation for a line, we need to know only two things, the slope and the y-intercept. We can figure the slope out by using "Rise over Run", which is where you take two points on a line, and measure how far the line goes up (rises) and how far the line goes to the right (runs) between the two points. Your points are (-4,-5) and (6,0), so the "Rise" would be how far the line goes up between 0 and -5, which is obviously 5. The "Run" would be how far the line goes to right between -4 and 6, which is 10. So using "Rise over Run", the slope would be 5/10, or 1/2. We now need to find the y-intercept. We can do this by using the point-slope equation. (Which is the equation for a line when you only know one point, and the slope): \[y-y _{1}=m(x-x _{1})\] where m is the slope and x and y with a little one under them are a specific point on the line. We know that m = 1/2 and we have the point (6,0) so we plug that into the equation to get \[y-0=(1/2)(x-6)\] which simplifies to \[y=(1/2)x -3\] And that is our equation for the line!

OpenStudy (anonymous):

thank you so much!